the leght of a rectangle is 5 cm greater than its width. the perimeter is 58 cm. find the dimension of the rectangle
Answer by checkley77(12844)

L=W+5
2L+2W=58
2(W+5)+2W=58
2W+10+2W=58
4W=58-10
4W=48
W=48/4
W=12 cm.
L=12+5=17 cm.
PROOF:
2*17+2*12=58
34+24=58
58=58